OpenStudy (anonymous):

Can somebody prove that 1,(9)=2?

OpenStudy (anonymous):

are you sure that is the question?

OpenStudy (anonymous):

sorry, I don't understand your question( It's just an interesting math problem, and the proof is quite pretty

OpenStudy (anonymous):

I don't know then...

OpenStudy (yanasidlinskiy):

Do you know what the answer is??

OpenStudy (anonymous):

yes)

OpenStudy (yanasidlinskiy):

Are you just trying to check if we know??? or what???lol:)

OpenStudy (anonymous):

Oh, I just think it's a nice problem)) when I found out the solution, I was surprised!

OpenStudy (yanasidlinskiy):

hmm....ok:) So whats the solution you found?!:)

OpenStudy (anonymous):

ok, I will write down.. 1,(9)=1,99999...(remind) let x=1,(9) then 10x=19,(9) 10x-x=19,(9)-1,(9)=18 9x=18 x=2 hence, x=1,(9)=2

OpenStudy (yanasidlinskiy):

That's nice!:) Bet no one else could figure that one out!

OpenStudy (anonymous):

hah, noo) it's a simple task for people, who are studying at university)

OpenStudy (yanasidlinskiy):

Cool. what grade are you in??

OpenStudy (anonymous):

I'm a freshman)

OpenStudy (yanasidlinskiy):

I'm actually suppose to be a freshman but since i took homeschool.....im now a...junior:) But you know........